OpenStudy (anonymous):

hiperbola: y^2=(x/b)^2+1 standard form????

OpenStudy (anonymous):

do you know the type of conic this equation gives?

OpenStudy (anonymous):

hyperbola

OpenStudy (anonymous):

right... so, what's the question?

OpenStudy (anonymous):

is it to write it in standard form?

OpenStudy (anonymous):

yes, i would like to know how do i write it in standard form

hartnn (hartnn):

the standard form is \(\huge \frac{y^2}{a^2}-\frac{x^2}{b^2}=1\) now can u write that in standard form ?

hartnn (hartnn):

just shift (x/b)^2 on left , by subtracting it from both sides.

hartnn (hartnn):

@cher u need further help?

OpenStudy (anonymous):

nope, thanks for the help

hartnn (hartnn):

welcome :)